OpenStudy (anonymous):

Please help and explain! I will fan and give medals! What is the solution for the equation 9^(2x) = (1/81)^(3+x) -2/3 -3/2 3 6

OpenStudy (anonymous):

start with the fact that \(\frac{1}{81}=9^{-2}\)

OpenStudy (anonymous):

so you have \[9^{2x}=9^{-2(3+x)}\] and then solve \[2x=-2(3+x)\]

OpenStudy (primeralph):